OpenStudy (anonymous):

The top of a ladder is 10 meters from the ground when the ladder leans against the wall at an angle of 35.5° with respect to the ground. If the ladder is moved by x meters toward the wall, it makes an angle of 54.5° with the ground, and its top is 14 meters above the ground. What is x rounded to the nearest meter?

OpenStudy (anonymous):



OpenStudy (anonymous):

@ABCKid333 do you know plane trigonometry?

OpenStudy (anonymous):

the problem can be solve using the cosine function. \[\cos \theta = \frac{ opposite side of given \angle }{ adjacent side ofgiven \angle }\]

OpenStudy (anonymous):

1st angle given is 35.5deg, and opposite side is 10m, therefore you can solve the adjacent side which is a+x in our drawing....

OpenStudy (anonymous):

2nd angle given is 54.5deg, opposite side to that is 14m, then we can solve "a" side...

OpenStudy (anonymous):

since we are looking for x meters, we can solve it by deducting "a" meters from "x+a" meters... that will give you the answer... :)